1) Which of the following ratios would be used to access a company’s ability to survive a four-year cyclical business downturn?

a.

Long-term debt to equity ratio

b.

Quick ratio

c.

Times interest earned

d.

Return on assets

2) Which of the following ratios would be best used to access the performance of company management in increasing shareholder wealth?

a.

Gross profit margin

b.

Economic Value Added

c.

Dividend yield

d.

Book value

3) What is the ABC Company’s coverage multiple (i.e., times interest earned ratio) given the following annual information: $1,200,000 of gross revenues, $300,000 cost of goods, $600,000 of selling, general and administrative expenses, $2,500,000 loan at 8% interest rate?

a.

4.5x

b.

3.0x

c.

1.5x

d.

6.0x
Answer #1

1. Which of the following ratios would be used to access a company’s ability to survive a four-year cyclical business downturn?

Answer : c : Times interest earned

explanation : In cyclical business downturn, earnings will go down, so most important aspect is, whether company is able to repay its debt holders or not. Because non-payment of interest can lead to liquidation of company.

2. Which of the following ratios would be best used to access the performance of company management in increasing shareholder wealth

Answer : b : Economic value added

explanation : Economic value added measures that after paying for all outsiders & after deducting cost of capital, what is left which is known as increase in shareholders wealth. If EVA is positive, we can say that value is generated by business for shareholders

All other ratios do not measure shareholder wealth

3. EBIT = sales - COGS - SG&D exp = 1200000 - 300000 - 600000 = 300000

interest = 2500000 x 8 % = 200000

coverage multiple = EBIT/interest = 300000/200000 = 1.5

Answer : c : 1.5x